OpenStudy (pphalke):

a basketball player scored 30 times during one game. she scored a total of 54 points, two for each two point shot and one for each free throw. how many two point shots did she make? how many free throws? so far i did x+y=30 2x+y=54 what are next steps?

OpenStudy (pphalke):

please help @ganeshie8 @ace-n-it

jimthompson5910 (jim_thompson5910):

x = # of two-point shots made y = # of free throws made solve the first equation for y x+y=30 x+y-x=30-x y = 30-x plug this into the other equation 2x+y=54 2x + (30-x) = 54 2x - x + 30 = 54 x+30 = 54 I'll let you finish up

OpenStudy (pphalke):

x=24

OpenStudy (pphalke):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me check @pphalke

jimthompson5910 (jim_thompson5910):

correct, nice job

jimthompson5910 (jim_thompson5910):

if x = 24, then what is y?

OpenStudy (pphalke):

do i put the x i got in 2x+y=54?

jimthompson5910 (jim_thompson5910):

you can plug it into that or into x+y = 30

jimthompson5910 (jim_thompson5910):

x+y = 30 may be easier actually y = 30-x may be easiest

OpenStudy (pphalke):

i got x=24 and y=6 when i pluged it in to 2x+y=54? does that sound right

OpenStudy (pphalke):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let's check x+y = 30 24+6 = 30 30 = 30 ... works 2x + y = 54 2*24 + 6 = 54 48 + 6 = 54 54 = 54 ... works answer is confirmed

OpenStudy (pphalke):

so 24 and 6 are right answers so that means the basketball player made 24 two point shots and 6 free throws

jimthompson5910 (jim_thompson5910):

bingo